Determine whether each of these functions is a bijection from R to R. a)f (x) = x^3
1
Expert's answer
2021-10-26T13:50:01-0400

Solution.

"f(x)=x^3,f:R\to R"



Let x1,x2 ∈R and let us assume f(x1)=f(x2).

So, "x_1^3=x_2^3\implies x_1=x_2."

Hence, we have f(x1)=f(x2)

​implies x1=x2.

So, f is one-one (injective).

Also we know


"-\infty<x<\infty\implies\newline
-\infty<x^3<\infty\implies\newline
-\infty<f(x)<\infty\implies\newline"


So, we clearly observe the Co-Domain is the same as the Range, so f(x) is surjective.

And hence f(x) is bijective.